# Effective Bertini theorems and zeros of $p$-adic forms of degree 7

### v1

@@ 4.2. Consequences
Here the superscript serves to remind us that $U^{(j)}\subset U$ has codimension $j$. To find $U^{(1)}$, we recognize that $\langle U,x\rangle\not\subset X_{f}$ for any $x\in(X_{f}-U)(\mathbb{F}_{q})$, since $\dim U$ is maximal, so applying Lemma 3.6 there are at most two possible codimension 1 $U^{(1)}\subset U$ defined over $\mathbb{F}_{q}$ for which $\langle U^{(1)},x\rangle\subset X_{f}$ for each $x\in(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})$. Since there are at most three $x\in(X_{f}\cap\langle u,v\rangle)(\mathbb{F}_{q})$ by Lemma 3.2, and more than 6 choices of $U^{(1)}$ by our hypotheses on $q$, there exists at least one satisfying (4.1). Repeating this argument, we find $U^{(j+1)}\subset U^{(j)}$ satisfying (4.1).
Suppose now that in $\langle U^{(1)},v,w\rangle\simeq\mathbb{P}^{s+1}$, $f$ vanishes on no linear subvarieties of dimension $s$. Then if $s>2$, the maximal dimension of such is s1s-1, and we may replace $U$ by $U^{(1)}$ and repeat the argument of this case. If $s=2$, then we have reduced to the case of $s=1$ above, so we are done.
Suppose to the contrary that for all of the $q+1$ codimension 1 $U_{i}^{(1)}\subset U$ defined over $\mathbb{F}_{q}$ satisfying $U^{(2)}\subset U_{i}^{(1)}$ (which implies (4.1) for $U_{i}^{(1)}$), we have that $f|_{\langle U_{i}^{(1)},v,w\rangle}$ vanishes on a linear subvariety $V_{i}$ of dimension $s$. Each $V_{i}$ meets the line $\langle v,w\rangle$ at exactly one of the at most three $\mathbb{F}_{q}$-points of $X_{f}\cap\langle v,w\rangle$. Thus whenever q>24q>{24}, there exist nine such $V_{i}$ containing the same $\mathbb{F}_{q}$-point of $X_{f}\cap\langle v,w\rangle$. Without loss of generality, assume $v\in V_{1},\ldots,V_{9}$.
@@
If $s>2$, let $V_{i}^{(3)}=V_{i}\cap U^{(2)}$, which has dimension s3s-3.333Here again the superscript serves to remind us that the codimension of $V_{i}^{(3)}\subset V_{i}$ is 3. Since $f$ vanishes on the span with $v$ of at most two codimension 1 linear subvarieties of $U^{(2)}$ by Lemma 3.6, at least five of the $V_{i}^{(3)}$ must coincide. Assume that for $V_{1},\ldots,V_{5}$ we have $V_{i}^{(3)}=V^{(3)}$. Now we have that $V^{(1)}=\langle V^{(3)},v,y\rangle\subset V_{i}$ for $i=1,\ldots,5$, so in fact the five $V_{i}$ meet in codimension 1. If $s=2$, then $V^{(3)}=\emptyset$ and $V^{(1)}=\langle v,y\rangle\subset V_{i}$ for all $i$, so we reach the same conclusion.
[…]
Our immediate goal is to use $V_{1}$, $V_{2}$, and $U$ to produce a plane $P$ which is skew to $V^{(1)}$ on which $f|_{P}\neq 0$ and vanishes on two lines. Then $P$ meets $V_{3},V_{4},V_{5}$ at three distinct points, and moreover these three points must not be colinear, lest $V_{3},V_{4},V_{5}$ span a space of dimension $s+1$, contradicting the discussion above. A final application of Lemma 3.5 then yields $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Let $v_{i}\in(V_{i}\cap U)-V^{(1)}$ for $i=1,2$. Thus we have $V_{i}\cap U=\langle V^{(1)}\cap U,v_{i}\rangle$ and $\langle v_{1},v_{2}\rangle\cap V^{(1)}=\emptyset$. Note that $\langle v_{1},v_{2}\rangle\subset X_{f}$, so applying Lemma 3.5 to $\langle V_{1},V_{2}\rangle$, we must have either that $f|_{\langle V_{1},V_{2}\rangle}$ vanishes on a smooth $\mathbb{F}_{q}$-point, vanishes on conjugate hyperplanes in $\langle V_{1},V_{2}\rangle$ that meet at a linear subvariety defined over $\mathbb{F}_{q}$ containing $\langle v_{1},v_{2}\rangle$, or vanishes on a third hyperplane $W\subset\langle V_{1},V_{2}\rangle$ containing $\langle v_{1},v_{2}\rangle$.
In the conjugate hyperplane case, we can find $\mathbb{F}_{q}$-points $u_{1}\in V_{1}-V^{(1)}$ and $u_{2}\in V_{2}-V^{(1)}$ not contained in the intersection of the conjugate hyperplanes. The line between them meets the conjugate hyperplanes at simple (geometric) points, so we are done.
Suppose now $W\subset\langle V_{1},V_{2}\rangle\cap X_{f}$ is defined over $\mathbb{F}_{q}$ with $\dim W=s$. Let $v_{2}^{\prime}\in(W-V^{(1)}-U)(\mathbb{F}_{q})$ so we have $\langle v_{1},v_{2}^{\prime}\rangle\subset X_{f}$ with the line meeting $U$ at $v_{1}$. By the maximality of $U$ and q>7q>{7}, successive applications of Lemma 3.6 show there exists $u^{\prime\prime}\in(U-\cup_{i=1}^{5}\langle V_{i},v_{1}\rangle)(\mathbb{F}_{q})$. Taking $P=\langle v_{1},v_{2}^{\prime},u^{\prime\prime}\rangle$, we see that $P\cap V^{(1)}=\emptyset$, while $f|_{P}$ vanishes on the lines $\langle v_{1},v_{2}^{\prime}\rangle\subset W$ and $\langle v_{1},u^{\prime\prime}\rangle\subset U$. By the discussion above, we have $X_{f}(F_{q})^{\mathrm{sm}}\neq\emptyset$, completing the proof. ∎
5. Proof of Theorem C
Proof of Theorem C.
@@ 5. Proof of Theorem C
Let $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ be a reduced form and denote by $f\in\mathbb{F}_{q}[x_{0},\ldots,x_{m}]$ its image under the residue map. If $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ then $X_{F}(K)\neq\emptyset$ by Hensel’s lemma.
@@
Assume $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}=\emptyset$. By Proposition 4.5, we find $u,v\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v\rangle\not\subset X_{f}$ and $f|_{\langle u,v\rangle}$ vanishes on a point of multiplicity 1, necessarily defined over $\overline{\mathbb{F}_{q}}$ but not $\mathbb{F}_{q}$.
@@
Possibly after a change of coordinates in $\operatorname{GL}_{n+1}(\mathcal{O}_{K})$, we may assume the $\overline{\mathbb{F}_{q}}$-points of $X_{f}$ along the line $\langle u,v\rangle$ are in the $x_{0}\neq 0$ affine patch. This allows us to find $(\mathbf{v}_{0},\mathbf{w}_{0})\in\mathbb{F}_{q}^{2n-1}$ such that
$f_{\mathbf{v}_{0},\mathbf{w}_{0},z}(X,0)=f(1,X+v_{1},w_{2}X+v_{2},\ldots,w_{n}X+v_{n})$
agrees with $f|_{\langle u,v\rangle}$. In particular, it vanishes on two distinct (singular) $\mathbb{F}_{q}$-points.
By Corollary 2.7, the hypothesis q>679q>{679} more than suffices to guarantee the existence of $\mathbf{z}_{0}$ such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}_{0}}(X,Y)$ has no geometric linear factors through the marked (geometric) point along the $Y=0$ line.
Re-homogenizing, $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}_{0}}(X,Y)$ corresponds to a degree 7 plane curve $C$ obtained by intersecting $X_{f}$ with a plane defined over $\mathbb{F}_{q}$. We have ensured that $C$ has a geometrically irreducible component defined over $\mathbb{F}_{q}$, namely that containing the $\overline{\mathbb{F}_{q}}$-point of multiplicity 1 on $\langle u,v\rangle$. Moreover, $C$ has at least two $\mathbb{F}_{q}$-points, $u$ and $v$.
@@
If $C$ is geometrically reducible, its geometrically irreducible factor has degree $d^{\prime}$ at most 5.444If $C$ has a factor of degree 6 then it also has a linear factor over $\mathbb{F}_{q}$, so we find a smooth point. Letting $C^{\prime}$ denote this irreducible factor, with genus $g^{\prime}$, we apply Lemma 3.3, finding that q>159q>{159} suffices to ensure $C(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ (the limiting case is $C^{\prime}$ a smooth quintic curve).
@@
It remains to consider the case of $C$ an irreducible degree 7 curve with at least two $\mathbb{F}_{q}$-points. If either of these points are smooth, we are done. If they are both singular, then the geometric genus is at most 13 (a smooth plane degree 7 plane curve has genus 15). Again applying (3.1), with $d^{\prime}=7$ and $g^{\prime}\leq 13$, we find $C(\mathbb{F}_{q})^{\mathrm{sm}}>0$ whenever q>679q>{679}.
@@
Under the hypotheses q>679q>{679} and $F$ reduced, we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, therefore $X_{F}(K)\neq\emptyset$. It is enough to check this only on reduced forms $F$ by Proposition 4.3(i), so we conclude $K$ is $C_{2}(7)$ when q>679q>{679}. ∎

### v2

@@ 4.2. Consequences
Here the superscript serves to remind us that $U^{(j)}\subset U$ has codimension $j$. To find $U^{(1)}$, we recognize that $\langle U,x\rangle\not\subset X_{f}$ for any $x\in(X_{f}-U)(\mathbb{F}_{q})$, since $\dim U$ is maximal, so applying Lemma 3.5 there are at most two possible codimension 1 $U^{(1)}\subset U$ defined over $\mathbb{F}_{q}$ for which $\langle U^{(1)},x\rangle\subset X_{f}$ for each $x\in(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})$. Since there are at most three $x\in(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})$ by Lemma 3.4, and more than 6 choices of $U^{(1)}$ by our hypotheses on $q$, there exists at least one satisfying (4.1). Repeating this argument, we find $U^{(j+1)}\subset U^{(j)}$ satisfying (4.1).
Suppose now that in $\langle U^{(1)},v,w\rangle\simeq\mathbb{P}^{s+1}$, $f$ vanishes on no linear subvarieties of dimension $s$. If $s=1$, we have reduced to the $s=0$ case above, so we are done. If $s>1$, we may replace $U$ by $U^{(1)}$ and repeat the argument of this case.
Suppose to the contrary that for all of the $q+1$ codimension 1 $U_{i}^{(1)}\subset U$ defined over $\mathbb{F}_{q}$ satisfying $U^{(2)}\subset U_{i}^{(1)}$ (which implies (4.1) for $U_{i}^{(1)}$), we have that $f|_{\langle U_{i}^{(1)},v,w\rangle}$ vanishes on a linear subvariety $V_{i}$ of dimension $s$. Each $V_{i}$ meets the line $\langle v,w\rangle$ at exactly one of the at most three $\mathbb{F}_{q}$-points of $X_{f}\cap\langle v,w\rangle$. Thus whenever q>18q>{18}, there exist five such $V_{i}$ containing the same $\mathbb{F}_{q}$-point of $X_{f}\cap\langle v,w\rangle$. Without loss of generality, assume $v\in V_{1},\ldots,V_{5}$.
@@
If $s>2$, let $V_{i}^{(3)}=V_{i}\cap U^{(2)}$, which has dimension s3s-3.333Here again the superscript serves to remind us that the codimension of $V_{i}^{(3)}\subset V_{i}$ is 3. Since $f$ vanishes on the span with $v$ of at most two codimension 1 linear subvarieties of $U^{(2)}$ by Lemma 3.5, at least three of the $V_{i}^{(3)}$ must coincide. Assume that for $V_{1},\ldots,V_{3}$ we have $V_{i}^{(3)}=V^{(3)}$. Now we have that $V^{(1)}=\langle V^{(3)},v,y\rangle\subset V_{i}$ for $i=1,\ldots,3$, so in fact the three $V_{i}$ meet in codimension 1. If $s=2$, then $V^{(3)}=\emptyset$ and $V^{(1)}=\langle v,y\rangle\subset V_{i}$ for all $i$, so we reach the same conclusion.
Suppose $\dim\langle V_{1},V_{2},V_{3}\rangle=s+1$. Then since $U\cap\langle V_{1},V_{2},V_{3}\rangle\neq\emptyset$, by Lemma 3.4(ii) we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. Thus we must have $\dim\langle V_{1},V_{2},V_{3}\rangle=s+2$, and we may apply Proposition 3.7. ∎
A similar approach gives a degree 11 analogue of Proposition 4.5.
Proposition 4.6.
Let $d=11$ and $n\geq 121$. Suppose q>7061q>{7061} and $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ is a reduced form of degree $11$. Then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $y,z\in X_{f}(\mathbb{F}_{q})$ such that $\langle y,z\rangle\not\subset X_{f}$ and $X_{f}\cap\langle y,z\rangle$ contains a geometric point of multiplicity 1.
Proof.
The proof follows the same lines as that of Proposition 4.5, using the intermediates from §3.2 in place of their analogues from §3.1.
The $s=0$ case follows from applying Proposition 4.3(iii) and Lemma 3.8 to show there exist noncolinear $u,v,w\in X_{f}(\mathbb{F}_{q})$, then invoking Lemma 3.10. We sketch the $s=1$ case below. We omit a justification of the $s\geq 2$ case, as it follows the same lines as the $s\geq 2$ of Proposition 4.5 with the modifications described herein.
Suppose $U\subset X_{f}$ is a line defined over $\mathbb{F}_{q}$ and $X_{f}$ contains no linear subvarieties of dimension 2 defined over $\mathbb{F}_{q}$. By Lemma 4.4, there exist $v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle v,w\rangle\not\subset X_{f}$ and $\langle v,w\rangle\cap U=\emptyset$. As in the proof of Proposition 4.5, if there exists $u\in U(\mathbb{F}_{q})$ for which $\langle u,v,w\rangle\cap X_{f}$ contains no lines defined over $F_{q}$, we apply Lemma 3.10 to obtain the desired result.
Assume instead for all $u\in U(\mathbb{F}_{q})$ there exists a line in $X_{f}\cap\langle u,v,w\rangle$ defined over $\mathbb{F}_{q}$. By Lemma 3.8(i), there are at most 20 such $u$ for which $u$ itself is contained in a line in $\langle u,v,w\rangle\cap X_{f}$. Since q>40q>{40}, there exist five points $u_{1},\ldots,u_{5}$ such that the associated lines $V_{1},\ldots,V_{5}$ all contain a common point. Note also that they are all skew to $U$.
Suppose $v\in V_{1},\ldots,V_{5}$ without loss of generality. If these five lines are coplanar, their span intersects $U$ at a point not contained in any $V_{i}$. Thus by Lemma 3.8(ii) we have $X_{f}(F_{q})^{\mathrm{sm}}=\emptyset$, since $f$ vanishes on no planes. If the five lines are not coplanar, any three of them which are not coplanar satisfy the hypotheses of Proposition 3.11, giving the desired conclusion. ∎
5. Proof of Theorem B
Proof of Theorem B.
@@ 5. Proof of Theorem B
Let $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ be a reduced form and denote by $f\in\mathbb{F}_{q}[x_{0},\ldots,x_{n}]$ its image under the residue map. If $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ then $X_{F}(K)\neq\emptyset$ by Hensel’s lemma. Assume $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}=\emptyset$.
Reducible case. Consider first that $f$ is geometrically reducible and write $f=\prod_{1\leq i\leq t}f_{i}^{e_{i}}$ over $\overline{\mathbb{F}_{q}}$. Since neither 7 nor 11 can be written as the sum of composites and $f$ has no geometric linear factors by Proposition 4.3(ii), at least one of the factors $f_{i}$ is defined over $\mathbb{F}_{q}$ with $e_{i}=1$.
By Theorem 2.2(i), there exists $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$ such that $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ corresponds to a form with the same factorization type as $f$ (see Example 2.5). Rehomogenizing, this corresponds to a plane curve $C\subset X_{f}$ with a geometrically integral component $C^{\prime}$ of multiplicity 1 defined over $\mathbb{F}_{q}$ of degree $d^{\prime}$ strictly less than 7 (resp. 11). Examining the possible factorizations of $f$, we always have such a component of degree $d^{\prime}\leq 3$ when $d=7$ (resp. $d^{\prime}\leq 7$ when $d=11$).
Applying Lemma 3.2 with $d^{\prime}\leq 3$ (resp. $d^{\prime}\leq 7$) we find that q>19q>{19} suffices (resp. q>942q>{942}) to guarantee $C(\mathbb{F}_{q})^{\mathrm{sm}}>0$. This is more than ensured by the hypotheses on $q$, so we conclude whenever $F$ is a reduced form with $f$ geometrically reducible, we have $X_{F}(K)\neq\emptyset$.
@@
Irreducible case. Assume $f$ is geometrically irreducible. By Proposition 4.5, we find $y,z\in X_{f}(\mathbb{F}_{q})$ such that $\langle y,z\rangle\not\subset X_{f}$ and $f|_{\langle y,z\rangle}$ vanishes on a point of multiplicity 1, necessarily defined over $\overline{\mathbb{F}_{q}}$ but not $\mathbb{F}_{q}$.
@@
Possibly after a change of coordinates in $\operatorname{GL}_{n+1}(\mathcal{O}_{K})$, we may assume the $\overline{\mathbb{F}_{q}}$-points of $X_{f}$ along the line $\langle y,z\rangle$ are in the $x_{0}\neq 0$ affine patch. This allows us to find $(\boldsymbol{\beta},\boldsymbol{\gamma})\in\mathbb{F}_{q}^{2n-1}$ such that
$f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,0)=f(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})$
vanishes on the same points as $f|_{\langle y,z\rangle}$. In particular, it vanishes on two distinct (singular) $\mathbb{F}_{q}$-points and has a (geometric) simple root $\alpha_{0}$ corresponding to the marked (geometric) point on $X_{f}$ along $\langle y,z\rangle$. We now treat the degree 7 and 11 cases separately.
(i) $d=7$. By Theorem 2.9, there is a polynomial $\Psi_{3}^{(\alpha_{0})}$ of degree at most 165 for which $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ has no nonconstant factor of degree at most 3 over $\overline{\mathbb{F}_{q}}$ vanishing on $(\alpha_{0},0)$ whenever $\Psi_{3}^{(\alpha_{0})}(\boldsymbol{\delta})\neq 0$. By the hypothesis on $q$, one such $\boldsymbol{\delta}$ must exist.
@@
Rehomogenizing, let $C\subset X_{f}$ denote the plane curve cut out by $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ for such a choice of $\boldsymbol{\delta}$. Let $C^{\prime}\subset C$ denote the geometrically integral component containing the point corresponding to $(\alpha_{0},0)$; we have $\deg C^{\prime}\in\{4,5,7\}$, [^4] If $C$ has a factor of degree 6 then it also has a linear factor over $\mathbb{F}_{q}$, so we find a smooth point. so it must appear with multiplicity 1 and be defined over $\mathbb{F}_{q}$. If $\deg C^{\prime}\neq 7$, then we apply Lemma 3.2 to find that q>159q>{159} suffices to ensure $C(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ (the limiting case here is $C^{\prime}$ a smooth quintic curve).
@@
It remains to consider the case of $C^{\prime}=C$ an irreducible degree 7 curve with at least two $\mathbb{F}_{q}$-points coming from $y$ and $z$. If either of these points are smooth, we are done. If they are both singular, then the geometric genus of $C$ is at most 13 (a smooth plane degree 7 plane curve has genus 15). Again applying (3.1), with $d^{\prime}=7$ and $g^{\prime}\leq 13$, we find $C(\mathbb{F}_{q})^{\mathrm{sm}}>0$ whenever q>679q>{679}.
@@
Under the hypotheses q>679q>{679} and $F$ reduced, we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, therefore $X_{F}(K)\neq\emptyset$. It is enough to check this only on reduced forms $F$ by Proposition 4.3(i), so we conclude $K$ is $C_{2}(7)$ when q>679q>{679}.
(ii) $d=11$. We follow the same approach. Theorem 2.9 gives $\Psi_{5}^{(\alpha_{0})}$ of degree at most 945 for which $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ has no nonconstant factor of degree at most 5 over $\overline{\mathbb{F}_{q}}$ vanishing on $(\alpha_{0},0)$ whenever $\Psi_{5}^{(\alpha_{0})}(\boldsymbol{\delta})\neq 0$.
As above, the assumptions on $q$ guarantee existence of one such $\boldsymbol{\delta}$, and $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ corresponds to a plane curve $C\subset X_{f}$ with at least two $\mathbb{F}_{q}$-points and a geometrically integral component $C^{\prime}$ of degree $d^{\prime}\in\{6,7,8,9,11\}$. If $d^{\prime}\neq 11$, q>3153q>{3153} suffices to guarantee a smooth point (the limiting case is $C^{\prime}$ a smooth curve of degree 9).
If $d^{\prime}=11$ we have $g^{\prime}\leq 43$ and applying (3.1) gives $C(\mathbb{F}_{q})^{\mathrm{sm}}>0$ whenever q>7393q>{7393}. Therefore $X_{F}(K)\neq\emptyset$ for all reduced $F$, and $K$ is $C_{2}(11)$ whenever q>7393q>{7393}. ∎